[INSTRM-2131] swap first two temp channels on b2, some pair on b4 Created: 16/Dec/23 Updated: 02/Feb/24 |
|
| Status: | Open |
| Project: | Instrument control development |
| Component/s: | ics_xcuActor |
| Affects Version/s: | None |
| Fix Version/s: | None |
| Type: | Task | Priority: | Normal |
| Reporter: | cloomis | Assignee: | cloomis |
| Resolution: | Unresolved | Votes: | 0 |
| Labels: | None | ||
| Remaining Estimate: | Not Specified | ||
| Time Spent: | Not Specified | ||
| Original Estimate: | Not Specified | ||
| Description |
|
Looks like the first two temp channels on b2 are swapped. There is also some permutation of the first three channels on b4, and/or maybe a dud/poorly bonded sensor: 2023-12-16T00:03:44.842 xcu_b1 : temps=150.6125,275.4944,142.5447,134.4267,277.4743,400.0000,400.0000,400.0000,400.0000,400.0000,162.2954,163.2170 2023-12-16T00:03:45.717 xcu_b2 : temps=276.9311,145.2434,152.1768,135.1048,278.7843,400.0000,400.0000,400.0000,400.0000,400.0000,162.8774,163.1753 2023-12-16T00:03:46.617 xcu_b3 : temps=154.7262,276.8176,152.7764,139.5380,278.6644,400.0000,400.0000,400.0000,400.0000,400.0000,158.2756,158.6268 2023-12-16T00:03:47.668 xcu_b4 : temps=148.3276,208.3578,275.9466,130.7394,277.4626,400.0000,400.0000,400.0000,400.0000,400.0000,162.2418,162.3472 Umm, the first three are the detector box, the Mangin, and the spider box. So I'd expect middle, high, low. Should look at rates to figure out b4. Swap back in the xcu actor I think. That knows which camera it belongs to, while the SMB program on the pi does not. But need to figure out the order on b4 first. |
| Comments |
| Comment by cloomis [ 02/Feb/24 ] |
|
Josh argues against this being a simple swap or two, given the tests done at JHU. Not sure what to do in that case, but I don't think it would be software work. We could plug test resistor packs in to see whether the temp board itself is behaving. |